# Gujarat-382355, India

### v1

@@ Abstract
A smoking gun for our current paradigm of the early universe would be direct evidence for the quantum mechanical origin of density perturbations which are conjectured to seed the large scale structure of our universe. A recently-proposed novel phenomenon is that of recoherence, wherein a specific interaction between the adiabatic and the entropic sector leads to the adiabatic mode retaining a coherent state after a transient increase in linear entropy. In this paper, we choose the most general Gaussian action and analyze the evolution of linear entropy, complexity of purification (COP), and complexity of formation (COF) to capture the interplay between decoherence and recoherence in this model. In the presence of two types of couplings that drive these two opposing characteristics, we highlight how COF is an efficient tool for diagnosing dynamics for such an open quantum system.
@@ 1 Introduction
The evolution of the Universe can be envisioned as a quantum circuit Bhattacharyya:2020rpy ; Bhattacharyya:2020kgu ; Haque:2021kdm ; Haque:2021hyw ; Haque:2024ldr ; Lehners:2020pem ; Liu:2021nzx ; Saha:2022onq ; Bhargava:2020fhl ; Li:2023ekd , where transitions between different phases of the evolution can be associated with the quantum complexity of each of these states. In this context, considering the primordial metric and matter fluctuations to be in a quantum state, the dynamics of complexity can provide valuable insights into the mechanisms of decoherence (and recoherence) of these degrees of freedom. One of the main motivations behind this lies in our lack of ability to track all the environment modes that cosmic observables interact with, and hence the requirement of considering the early universe as an open quantum system. Circuit complexity NL1 ; NL2 ; NL3  [^2] Interested readers are referred to some of the works done in the context of Nielsen complexity NL1 ; NL2 ; NL3 for quantum field theory and quantum many-body systems Jefferson ; Chapman:2017rqy ; Bhattacharyya:2018wym ; Caputa:2017yrh ; Ali:2018fcz ; Bhattacharyya:2018bbv ; Hackl:2018ptj ; Khan:2018rzm ; Camargo:2018eof ; Ali:2018aon ; Caputa:2018kdj ; Guo:2018kzl ; Bhattacharyya:2019kvj ; Flory:2020eot ; Erdmenger:2020sup ; Ali:2019zcj ; Bhattacharyya:2019txx ; Caceres:2019pgf ; Bhattacharyya:2020art ; Liu_2020 ; Susskind:2020gnl ; Chen:2020nlj ; Czech:2017ryf ; Chapman:2018hou ; Geng:2019yxo ; Guo:2020dsi ; Couch:2021wsm ; Erdmenger:2021wzc ; Chagnet:2021uvi ; Koch:2021tvp ; Banerjee:2022ime ; Bhattacharyya:2022ren ; Bhattacharyya:2023sjr ; Bhattacharyya:2022rhm ; Bhattacharyya:2021fii ; Bhattacharyya:2020iic ; ShashiPRB ; casagrande2023complexity ; Craps:2023rur ; Jaiswal:2021tnt ; Jaiswal:2020snm ; Haque:2021kdm ; Bhattacharya:2022wlp . This list is by no means exhaustive and the reader is referred to the following reviews and thesis Chapman:2021jbh ; Bhattacharyya:2021cwf ; Katoch:2023etn ; Aguilar-Gutierrez:2024yzu , and references therein, for more details. in open quantum systems has been extensively explored across various models Bhattacharyya:2021fii ; Bhattacharyya_2024 ; Bhattacharyya:2022rhm to understand its significance in quantifying the dynamics of the reduced density matrix corresponding to the system modes. Of particular interest is the study of decoherence in the flat-space Caldeira-Leggett model PhysRevLett.46.211 , as explored in Bhattacharyya:2022rhm , where the time evolution of the complexity of purification (COP) Ali:2018fcz was compared to other standard decoherence detection tools, such as linear entropy and entanglement negativity. Interestingly, this analysis could be successfully extended to curved spacetime, specifically for the two-field model in de Sitter background Bhattacharyya_2024 . For both flat and curved spacetime scenarios, circuit complexity demonstrated significant sensitivity to the decoherence mechanism. In light of this finding – sensitivity of quantum complexity as a tool for decoherence – it is natural to ask if complexity can be a quantifier for more exotic phenomena such as the recoherence mechanism established in Colas:2022kfu .
@@
This text is organized as follows: In section 2, we introduce the toy model of two interacting fields. In section 3 and section 4, we briefly review linear entropy and complexity measures, respectively. In section 5, we discuss the time evolution of linear entropy and complexity measures. Finally, we conclude with a discussion and outline potential future directions.
@@ 2 The Model
The motivation for considering this toy model arises from identifying the $\varphi$ field with the adiabatic fluctuation while $\chi$ plays the role of fluctuations in one of the many heavy fields that are typically present during inflation. Indeed, from an EFT point of view, the leading order mixing term between these fields would be expressed as given above. When the inflaton fluctuations are massless (a statement that is true for the primordial curvature perturbation $\zeta$ to leading order in slow-roll), then the $\lambda$ term effectively switches off, a well-known fact for the EFT of adiabatic mode (see, for instance, Colas:2022kfu ; Assassi:2013gxa ). Our Lagrangian matches theirs on considering the field redefinition $\varphi\equiv\sqrt{2\epsilon}M_{\text{Pl}}\,\zeta$. However, we keep both these terms to give the most general two-field model up to Gaussian couplings and this versatility would be useful in exploring both mixedness of the state, and its ability to retain coherence, by tuning these parameters. More explicitly, note that by turning off the momentum-field interaction (i.e., setting $\rho\rightarrow 0$), the model reduces to the so-called curved-space Caldeira-Leggett framework, which has been extensively studied recently in Colas2022 for demonstrating quantum decoherence. In Bhattacharyya_2024 , this approach was extended to demonstrate that quantum complexity can also serve as a useful tool for detecting decoherence.
@@
$\Sigma^{(\varphi\varphi)}=\begin{pmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{pmatrix}\,.$ (12)
Since we are interested in quantum informatic properties of the system field, as a response to tracing out the heavy entropic mode, this $\Sigma^{(\varphi\varphi)}$ block of the covariance matrix will be useful for computing both the linear entropy and complexity measures. In the following sections, we provide a brief review of these quantities, followed by an analysis of their time evolution to study decoherence and recoherence.
@@ 3 Linear Entropy
$S_{l}=1-\frac{1}{4\det(\Sigma^{(\varphi\varphi)})}\,.$ (14)
@@ 4 Complexity Measures: COP and COF
where $\omega$ and $\omega_{0}$ are the frequencies of the target state and reference state respectively. However, the method described above is for a pure state by construction. Extending the computation of complexity to open quantum systems is still an active area of research, however the complexity of purification (COP) Caceres:2019pgf ; Agon:2018zso ; Camargo:2018eof ; Ghodrati:2019hnn ; Bhattacharyya:2020iic is one such extension. What follows is a description of how COP is calculated.
@@
Given that the choice of ancillary Hilbert space is not unique (and hence the purification process is dependent on this choice), a set of pure states |$\ket{\Psi}_{\alpha,\beta,\gamma,...}$ can be constructed, parametrised by $\alpha,\beta,\gamma,...$ which satisfy the purification requirement. We then minimize the quantity of interest (in this case, the complexity $\mathcal{C}$) with respect to these parameters. Thus, the COP is obtained as:
@@
where |$\ket{\Psi}_{R}$ is the reference state and $\mathcal{C}$ is the complexity as calculated in the enlarged Hilbert space. Given that the purified wavefunction has a Gaussian structure in our model, it can be parametrized as
@@
where $\rm{z}$ and $\rm{z_{anc}}$ belong to the Hilbert spaces $\mathcal{H}$ and $\mathcal{H}_{\rm anc}$ respectively. The parameters $\alpha,\beta$ and $\tau$ can be specified in the following way: trace out the ancillary degrees of freedom of the density matrix |$\ket{\psi}\bra{\psi}$ | to find the reduced density matrix $\hat{\rho}_{\rm{mix}}$ which also has a Gaussian structure and hence can readily be linked to the covariance matrix $\Sigma^{(\varphi\varphi)}$.
Thus, the COP for our model can be computed as Bhattacharyya:2021fii :
@@
The frequency $\omega_{0}$ of the reference state is set to 1 for convenience. Note that no direct expression is provided for $\text{Im}(\beta)$. Thus $\text{Im}(\beta)$ is treated as a free parameter, and in order to compute COP, the expression in (20) is minimised over $\text{Im}(\beta)$.
@@ 5.1 Field-Field Coupling $\lambda$ vs Momentum-Field Coupling $\rho$
Figure 1: Time evolution of linear entropy $(S_{l})$ in the special case where either the field-field coupling (fig. 1(a)) or the momentum-field coupling (fig. 1(b)) is enabled. Here, $H=1$, $M=4$ and $k=1$ in both plots.
We immediately notice that the late-time behavior of the linear entropy is qualitatively different for these two cases. In the case with only the field-field coupling (fig. 1(a)), the linear entropy follows a sigmoid curve, suddenly increasing at a timescale inversely proportional to the field-field coupling $\lambda$ and saturates near 1. This suggests that at late times the system has decohered completely. On the other hand, in the case with only the momentum-field coupling (fig. 1(b)), the linear entropy undergoes a transient period of increase after which it decreases and settles back near 0 (on choosing a region of parameter space such that the entropic field is sufficiently heavy, which is the case for all the cases considered in fig. 1(b)). This suggests that the system transiently becomes more mixed, and then recoheres to remain pure at late times. The timescale at which the system starts to recohere is independent of the momentum-field coupling strength $\rho$. It is important to note that while the degree to which decoherence takes before recoherence increases with $\rho$ for the momentum-field coupling, it certainly does not lead to complete decoherence as observed in the field-field coupling case. This feature highlights the role of the specific interaction terms: the momentum-field coupling causes partial decoherence before the system recoheres, in contrast to the case of the field-field coupling which induces complete decoherence.
@@
Figure 2: Time evolution of complexity of purification (COP) in the special case where either the field-field coupling (fig. 2(a)) or the momentum-field coupling(fig. 2(b)) is enabled. Here, $H=1$, $M=4$ and $k=1$ in both plots.
Now we will compare the above behavior with that of the measures of complexity, namely COP and COF. With only the field-field coupling enabled (fig. 2(a)), the COP is initially fairly close to 0, followed by a short period of intermediate growth, and finally a late-time rapid and unbounded growth. The intermediate growth and late-time growth periods are separated by a sudden bump. The late-time growth starts at a timescale similar to the saturation timescale of linear entropy. Moreover, the timescale of this COP growth is inversely proportional to the coupling strength $\lambda$. This is indicative of the fact that COP is sensitive to decoherence for the case of field-field coupling (fig. 2(a)) Bhattacharyya_2024 . The time evolution of COF also shows a similar feature as the late-time growth of COP (see fig. 3(a)) for the field-field coupling, hence a similar correspondence can be suggested between the COF growth and the decoherence of the system.
Interestingly, when only the momentum-field coupling is enabled, the time evolution of COP does not show any clear feature that is indicative of recoherence (see fig. 2(b)). We suspect that this is due to the large squeezing effect as mentioned in the earlier section. The background squeezing obfuscates any signature of the decrasing linear entropy in the COP and one needs to separate the two to find any discernable signal. In this regard, COF will be a useful tool to probe the recoherence of the system. As demonstrated in fig. 3(b), COF shows a sudden increase followed by a rapid decrease during the transient period of decoherence to recoherence, and finally saturates rather than linear growth.
This saturation is a distinctive feature of COF and to understand its implication we will compare its timescale with that of the features observed in the time evolution of linear entropy. The time evolution of linear entropy shown in fig. 1(b) suggests that the saturation of COF signifies the completion of recoherence of the system. In addition, we notice that the saturation value of COF increases with the coupling strength $\rho$. This is similar to the case of linear entropy where the peak before recoherence increases with increasing coupling strength.
@@
(a) Field-field coupling
(b) Momentum-field coupling
Figure 3: Time evolution of complexity of formation (COF) in the special case where either the field-field coupling (fig. 3(a)) or the momentum-field coupling (fig. 3(b)) is enabled. Here, $H=1$, $M=4$ and $k=1$ in both plots.

### v2

@@ Abstract
A smoking gun for our current paradigm of the early universe would be direct evidence for the quantum mechanical origin of density perturbations which are conjectured to seed the large scale structure of our universe. A recently-proposed novel phenomenon is that of recoherence, wherein a specific interaction between the adiabatic and the entropic sector leads to the adiabatic mode retaining a coherent state after a transient increase in linear entropy. In this paper, we choose the Gaussian action allowing for both field-field and field-momentum coupling, and analyze the evolution of linear entropy, complexity of purification (COP), and complexity of formation (COF) to capture the interplay between decoherence and recoherence in this model. In the presence of these two types of couplings that drive these two opposing characteristics, we highlight how COF is an efficient tool for diagnosing dynamics for such an open quantum system.
@@ 1 Introduction
The evolution of the Universe can be envisioned as a quantum circuit Bhattacharyya:2020rpy ; Bhattacharyya:2020kgu ; Haque:2021kdm ; Haque:2021hyw ; Lehners:2020pem ; Liu:2021nzx ; Saha:2022onq ; Li:2023ekd , where transitions between different phases of the evolution can be associated with the quantum complexity of each of these states. In this context, considering the primordial metric and matter fluctuations to be in a quantum state, the dynamics of complexity can provide valuable insights into the mechanisms of decoherence (and recoherence) of these degrees of freedom. One of the main motivations behind this lies in our lack of ability to track all the environment modes that cosmic observables interact with, and hence the requirement of considering the early universe as an open quantum system. Circuit complexity NL1 ; NL2 ; NL3  [^2] Interested readers are referred to some of the works done in the context of Nielsen complexity NL1 ; NL2 ; NL3 for quantum field theory and quantum many-body systems Jefferson ; Chapman:2017rqy ; Bhattacharyya:2018wym ; Caputa:2017yrh ; Ali:2018fcz ; Bhattacharyya:2018bbv ; Hackl:2018ptj ; Khan:2018rzm ; Camargo:2018eof ; Ali:2018aon ; Caputa:2018kdj ; Guo:2018kzl ; Bhattacharyya:2019kvj ; Flory:2020eot ; Erdmenger:2020sup ; Ali:2019zcj ; Bhattacharyya:2019txx ; Caceres:2019pgf ; Bhattacharyya:2020art ; Liu_2020 ; Susskind:2020gnl ; Chen:2020nlj ; Czech:2017ryf ; Chapman:2018hou ; Geng:2019yxo ; Guo:2020dsi ; Couch:2021wsm ; Erdmenger:2021wzc ; Chagnet:2021uvi ; Koch:2021tvp ; Banerjee:2022ime ; Bhattacharyya:2022ren ; Bhattacharyya:2023sjr ; Bhattacharyya:2022rhm ; Bhattacharyya:2021fii ; Bhattacharyya:2020iic ; Craps:2023rur ; Jaiswal:2021tnt ; Jaiswal:2020snm ; Haque:2021kdm ; Bhattacharya:2022wlp . This list is by no means exhaustive and the reader is referred to the following reviews and thesis Chapman:2021jbh ; Bhattacharyya:2021cwf ; Katoch:2023etn ; Aguilar-Gutierrez:2024yzu , and references therein, for more details. in open quantum systems has been extensively explored across various models Bhattacharyya:2021fii ; Bhattacharyya_2024 ; Bhattacharyya:2022rhm to understand its significance in quantifying the dynamics of the reduced density matrix corresponding to the system modes. Of particular interest is the study of decoherence in the flat-space Caldeira-Leggett model PhysRevLett.46.211 , as explored in Bhattacharyya:2022rhm , where the time evolution of the complexity of purification (COP) Ali:2018fcz was compared to other standard decoherence detection tools, such as linear entropy and entanglement negativity  [^3] For other measures of decoherence in the context of cosmology interested readers are referred to some of theses papers Martin:2021znx ; Martin:2022kph  .. Interestingly, this analysis could be successfully extended to curved spacetime, specifically for the two-field model in de Sitter background Bhattacharyya_2024 . For both flat and curved spacetime scenarios, circuit complexity demonstrated significant sensitivity to the decoherence mechanism. In light of this finding – sensitivity of quantum complexity as a tool for decoherence – it is natural to ask if complexity can be a quantifier for more exotic phenomena such as the recoherence mechanism established in Colas:2022kfu .
@@
This text is organised as follows: In section 2, we introduce the toy model of two interacting fields. In section 3 and section 4, we briefly review linear entropy and complexity measures, respectively. In section 5, we discuss the time evolution of linear entropy and complexity measures. Finally, we conclude with a discussion and outline potential future directions.
@@ 2 The Model
The motivation for considering this toy model arises from identifying the $\varphi$ field with the adiabatic fluctuation while $\chi$ plays the role of fluctuations in one of the many heavy fields that are typically present during inflation. Indeed, from an EFT point of view, the leading order mixing term between these fields would be expressed as given above. When the inflaton fluctuations are massless (a statement that is true for the primordial curvature perturbation $\zeta$ to leading order in slow-roll), then the $\lambda$ term effectively switches off, a well-known fact for the EFT of adiabatic mode (see, for instance, Colas:2022kfu ; Assassi:2013gxa ). Our Lagrangian matches theirs on considering the field redefinition $\varphi\equiv\sqrt{2\epsilon}M_{\text{Pl}}\,\zeta$. However, we keep both these terms to allow for a more general two-field model, restricted to Gaussian couplings, that is interesting from a phenomenological point of view. As it turns out, this versatility would be useful in exploring both mixedness of the state, and its ability to retain coherence, by tuning these parameters. More explicitly, note that by turning off the momentum-field interaction (i.e., setting $\rho\rightarrow 0$), the model reduces to the so-called curved-space Caldeira-Leggett framework, which has been extensively studied recently in Colas2022 for demonstrating quantum decoherence. In Bhattacharyya_2024 , this approach was extended to demonstrate that quantum complexity can also serve as a useful tool for detecting decoherence.
@@
$\Sigma^{(\varphi\varphi)}=\begin{pmatrix}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{21}&\Sigma_{22}\end{pmatrix}\,,$ (12)
where $\Sigma_{21}=\Sigma_{12}$ as follows from eq. 9, leaving $\Sigma^{\left(\varphi\varphi\right)}$ with only three independent entries. Since we are interested in quantum informatic properties of the system field, as a response to tracing out the heavy entropic mode, this $\Sigma^{(\varphi\varphi)}$ block of the covariance matrix will be useful for computing both the linear entropy and complexity measures. In the following sections, we provide a brief review of these quantities, followed by an analysis of their time evolution to study decoherence and recoherence.
@@ 3 Linear Entropy
$S_{l}=1-\frac{1}{4\det(\Sigma^{(\varphi\varphi)})}\,,$ (14)
a result that holds for a Gaussian state.
@@ 4 Complexity Measures: COP and COF
where $\omega$ and $\omega_{0}$ are the frequencies of the target state and reference state respectively. The reference state wavefunction is of the form $\psi_{R}(t=0,x)\sim e^{-\frac{\omega_{0}\,x^{2}}{2}}$ and $\psi_{T}(t,x)\sim e^{-\frac{\omega(t)\,x^{2}}{2}}\,,$ where $\omega(t)$ is in general complex and $\omega_{0}$ is real. $\psi_{T}(t,x)$ is obtained by time-evolving $\psi_{R}(t=0,x)$ with the harmonic oscillator Hamiltonian. Also, the fundamental block for the circuit for this case is the gates generated by exponentiating scaling operator $\frac{i}{2}(\hat{x}\hat{p}+\hat{p}\hat{x})\,,$ where the hat symbol is used to denote the operators and $[\hat{x},\hat{p}]=i\,.$ For more details readers are referred to Sec. (3.2) of Ali:2018fcz . Furthermore, one can easily check that, when $\omega\rightarrow\omega_{0},$ the complexity defined in eq. (15), vanishes as the target and reference coincide in this limit.
However, the method described above is for a pure state by construction. Extending the computation of complexity to open quantum systems is still an active area of research, however the complexity of purification (COP) Caceres:2019pgf ; Agon:2018zso ; Camargo:2018eof ; Ghodrati:2019hnn ; Bhattacharyya:2020iic is one such extension. What follows is a description of how COP is calculated.
@@
Given that the choice of ancillary Hilbert space is not unique (and hence the purification process is dependent on this choice), a set of pure states |$\ket{\Psi}_{\alpha,\beta,\gamma,...}$ can be constructed, parametrised by $\alpha,\beta,\gamma,...$ which satisfy the purification requirement. We then minimize the quantity of interest (in this case, the complexity $\mathcal{C}$) with respect to these parameters. Thus, the COP is obtained as Bhattacharyya:2020iic :
@@
where |$\ket{\Psi}_{R}$ is the reference state and $\mathcal{C}$ is the complexity as calculated in the enlarged Hilbert space. Following Caceres:2019pgf ; Bhattacharyya:2018sbw ; Bhattacharyya:2020iic , we choose to use minimal purification such that the size system is same as the ancillary system as well as restrict ourselves to Gaussian purification [^4] It will be interesting to go beyond the minimal purification scheme. We leave this for future studies.. Given these, the purified state can be parametrized as
@@
where $\rm{z}$ and $\rm{z_{anc}}$ belong to the Hilbert spaces $\mathcal{H}$ and $\mathcal{H}_{\rm anc}$ respectively. The parameters $\alpha,\beta$ and $\tau$ can be specified in the following way: trace out the ancillary degrees of freedom of the density matrix |$\ket{\psi}\bra{\psi}$ | to find the reduced density matrix $\hat{\rho}_{\rm{mix}}$ which also has a Gaussian structure and hence can readily be linked to the covariance matrix $\Sigma^{(\varphi\varphi)}$. Then following Bhattacharyya:2020iic , this fixes $\alpha,\tau$ and $\text{Re}(\beta)\,.$
Now notice that the form of $\psi(\rm{z},\rm{z}_{\rm{anc}})$ is that of two-coupled oscillators. Then, a straightforward generalization of eq. (15) for the two-oscillator case [^5] Readers are referred to Ali:2018fcz for more details. Ali:2018fcz contains an expression for the circuit complexity for N-coupled oscillators in terms of the normal-mode frequencies. The expression under the square root in eq. (20) is just a special case of the expression quoted in eq. (3.44) of Ali:2018fcz . gives an expression for the COP for our model as Bhattacharyya:2020iic :
@@
The frequency $\omega_{0}$ of the reference state is set to 1 for convenience. $\omega_{1}$ and $\omega_{2}$ are the normal-mode frequencies in the purified space. Note that eq. (23) fixes all the parameters except $\text{Im}(\beta)$ for a given $\Sigma^{(\varphi\varphi)}\,.$ Thus $\text{Im}(\beta)$ is treated as a free parameter, and in order to compute COP, the expression in (20) is minimised over $\text{Im}(\beta)$.
@@ 5.1 Field-Field Coupling $\lambda$ vs Momentum-Field Coupling $\rho$
Figure 1: Time evolution of linear entropy $(S_{l})$ in the special case where either the field-field coupling (Fig. (1(a))) or the momentum-field coupling (Fig. (1(b))) is enabled. Here, $H=1$, $M=4$ and $k=1$ in both plots.
We immediately notice that the late-time behavior of the linear entropy is qualitatively different for these two cases. In the case with only the field-field coupling (Fig. (1(a))), the linear entropy follows a sigmoid curve, suddenly increasing at a timescale inversely proportional to the field-field coupling $\lambda$ and saturates near 1. This suggests that at late times the system has decohered completely. On the other hand, in the case with only the momentum-field coupling (Fig. (1(b))), the linear entropy undergoes a transient period of increase after which it decreases and settles back near 0 (on choosing a region of parameter space such that the entropic field is sufficiently heavy, which is the case for all the cases considered in Fig. (1(b))). This suggests that the system transiently becomes more mixed, and then recoheres to remain pure at late times. The timescale at which the system starts to recohere is independent of the momentum-field coupling strength $\rho$. It is important to note that while the degree to which decoherence takes before recoherence increases with $\rho$ for the momentum-field coupling, it certainly does not lead to complete decoherence as observed in the field-field coupling case. This feature highlights the role of the specific interaction terms: the momentum-field coupling causes partial decoherence before the system recoheres, in contrast to the case of the field-field coupling which induces complete decoherence.
@@
Figure 2: Time evolution of complexity of purification (COP) in the special case where either the field-field coupling (Fig. (2(a))) or the momentum-field coupling(Fig. (2(b))) is enabled. Here, $H=1$, $M=4$ and $k=1$ in both plots.
Now we will compare the above behavior with that of the measures of complexity, namely COP and COF. With only the field-field coupling enabled (Fig. 2(a)), the COP is initially fairly close to 0, followed by a short period of intermediate growth, and finally a late-time rapid and unbounded growth. The intermediate growth and late-time growth periods are separated by a sudden bump. The late-time growth starts at a timescale similar to the saturation timescale of linear entropy. Moreover, the timescale of this COP growth is inversely proportional to the coupling strength $\lambda$. This is indicative of the fact that COP is sensitive to decoherence for the case of field-field coupling (Fig. (2(a))) Bhattacharyya_2024 . The time evolution of COF also shows a similar feature as the late-time growth of COP (see Fig. (3(a))) for the field-field coupling, hence a similar correspondence can be suggested between the COF growth and the decoherence of the system.
Interestingly, when only the momentum-field coupling is enabled, the time evolution of COP does not show any clear feature that is indicative of recoherence (see Fig. (2(b))). We suspect that this is due to the large squeezing effect as mentioned in the earlier section. The background squeezing obfuscates any signature of the decrasing linear entropy in the COP and one needs to separate the two to find any discernable signal. In this regard, COF will be a useful tool to probe the recoherence of the system. As demonstrated in Fig. (3(b)), COF shows a sudden increase followed by a rapid decrease during the transient period of decoherence to recoherence, and finally saturates rather than linear growth.
This saturation is a distinctive feature of COF and to understand its implication we will compare its timescale with that of the features observed in the time evolution of linear entropy. The time evolution of linear entropy shown in Fig. (1(b)) suggests that the saturation of COF signifies the completion of recoherence of the system. In addition, we notice that the saturation value of COF increases with the coupling strength $\rho$. This is similar to the case of linear entropy where the peak before recoherence increases with increasing coupling strength.
(a) Field-field coupling
(b) Momentum-field coupling
Figure 3: Time evolution of complexity of formation (COF) in the special case where either the field-field coupling (Fig. (3(a))) or the momentum-field coupling (Fig. (3(b))) is enabled. Here, $H=1$, $M=4$ and $k=1$ in both plots.